OpenStudy (anonymous):

Can someone check my answers please? A. solve each porportion 1. 3/4 = ?/12 = 9 2. 4/? = 12/18 = 6 3. 30/18 = 10/? = 6 4. 7/14 = ?/4 = 2 5. 20/? = 4/11 =55 6. ?/30 = 10/12 =25 7. ?/42 = 10/12 =35 8. 5/24 = 4/? =19.2 9. 5/8 = ?/40 = 25 B. Use porportions to solve these problems 10. Toya gets paid an hourly wage. During a 40- hour week, Toya earns 4240 in her job as a receptionist. This week she worked only 35 hours. How much did she earn? 40/240 = 35/? =210 11. Kelsey swims 35 laps in 25 mintues. How many laps does he swim in 15 mi

OpenStudy (anonymous):

11. Kelsey swims 35 laps in 25 mintues. How many laps does he swim in 15 minutes? 35/25 = ?/15 = 21 12. Sarah is a photographer. For exery 15 rolls of fil she shoots, 5 are black and white. if Sara uses 51 rolls of film, how many rolls are in black and white? 5/15 = ?/51 =15 13. Dixie walks 45 minutes a day. how many hours will she walk in 8days? 45/1 = ?/8 =360

OpenStudy (anonymous):

14. one meter is 39.37 inches. how many inches is 2.5 meters 1/39.37 = 2.5/? =98.425 inches 15. noriko is in charge of the airpot parkin lots. Lot A has 4000 spaces. Today there are 2500 cars in lot A. Lot B has 2000 spaces. If both lots have the same ratio of cars to spaces, how many cars are in parking lot B? 4000/2500 = 2000/? =1250

OpenStudy (anonymous):

thats the second part that got cut off..

OpenStudy (anonymous):

all of it seems correct except recheck question 10.

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

ok thank!